# On multi-graded Proj schemes

### v3

@@ 3.5 Proj, blowups and dilatations
In the setting above, the scheme ${\operatorname{Proj}}^{\mathbb{Z}_{I}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)$ is covered by the open subschemes ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])$ where $(a_{i}:i\in I)$ runs over families such $a_{i}\in J_{i}$ for all $i\in I$.
@@
By Corollary 3.10, the lemma will follow if we prove that any relevant element in the $\mathbb{Z}_{I}$-graded ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$ belongs to the sum of the ideals $({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)\cdot\prod_{i}a_{i}$ where $(a_{i}:i\in I)$ runs over the families as above. Now it is clear that this sum is the sum of the graded components in ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$ associated with elements in $\mathbb{Z}_{I}$ all of whose components are positive. The claim therefore follows from the observation that the degree of any relevant element in ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$ must have positive components, because if the $i$-component of the degree of a homogeneous element is 00, then the same will hold for all divisors of all of its powers, so that this element cannot be relevant. ∎
We now work in a relative setting; we therefore fix a base $X$, and let $\{Y_{i}:i\in I\}$ be a finite family of closed subschemes.
@@
Clearly, a blowup of $X$ with center $\{Y_{i}:i\in I\}$ is unique up to unique isomorphism if it exists. For each $i$, let $\mathcal{J}_{i}\subset{\mathcal{O}}_{X}$ be the quasi-coherent ideal of $Y_{i}$, so that $Y_{i}=V(\mathcal{J}_{i})$. Let
$\mathrm{Bl}_{\{\mathcal{J}_{i}:i\in I\}}\mathcal{O}_{X}=\bigoplus_{\nu\in\mathbb{N}_{I}}\mathcal{J}^{\nu}$
be the associated Rees algebra (cf. e.g. [Ma23b, §3.1]). This is canonically a $\mathbb{Z}_{I}$-graded quasi-coherent $\mathcal{O}_{X}$-algebra. Using the construction of §3.4 we now define an $X$-scheme
$\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X:={\operatorname{Proj}}^{\mathbb{Z}_{I}}_{X}(\mathrm{Bl}_{\{\mathcal{J}_{i}:i\in I\}}\mathcal{O}_{X}).$
We will prove in Proposition 3.30 below that this scheme is a blowup of $X$ with center $\{Y_{i}:i\in I\}$ in the sense of Definition 3.25. (The case $|I|=1$ of this statement is classical, see e.g. [StP, Tag 01OF].)
@@
Since the property of being separated is local on the target (see e.g. [StP, Tag 02KU]), we can assume $X={\operatorname{Spec}}(A)$ is affine, and by [StP, Tag 01KV] the claim will follow if in this case we prove that $\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$ is a separated scheme. We proceed with usual local notations from the beginning of the subsection. Let ${\mathcal{F}}$ be the collection of relevant homogeneous families $(a_{i}:i\in I)$ with $a_{i}\in J_{i}$ for all $i\in I$. By Lemma 3.24, ${\operatorname{Proj}}^{M}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)={\operatorname{Proj}}^{M}_{{\mathcal{F}}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)$. For any family $T=(a_{i}:i\in I)$ in ${\mathcal{F}}$, we have $C_{T}=(\mathbb{R}_{\geqslant 0})_{I}\subset\mathbb{R}_{I}$. So if $T,T^{\prime}$ belong to ${\mathcal{F}}$, $C_{T}\cap C_{T^{\prime}}$ has nonempty interior. Now Proposition 3.6 finishes the proof. ∎
@@
This statement is local on $X$, so we can assume $X={\operatorname{Spec}}(A)$ for some ring $A$, and write $J_{i}$ for the global sections of ${\mathcal{J}}_{i}$. In view of Lemma 3.24, in this case it suffices to prove that for any family $(a_{i}:i\in I)$ with $a_{i}\in J_{i}$ for any $i$, if $C$ is an effective Cartier divisor on $X$ then ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])\times_{X}C$ is an effective Cartier divisor on ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])$. This statement is a special case of [Ma23b, Proposition 3.18]. ∎
In the following statement we use the notation of dilatations of schemes from [Ma23b] (see also [MRR23, DMdS23]). For simplicity of notation we set
$X^{\prime}:=\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X,$
and for $i\in I$ we set
$Y_{i}^{\prime}:=Y_{i}\times_{X}X^{\prime}.$
Proposition 3.28 ((Compatibility)).
Let $\{D_{i}:i\in I\}$ be a family of locally principal closed subschemes of $X$, such that $Y_{i}\subset D_{i}$ for all $i$, and set $D_{i}^{\prime}:=D_{i}\times_{X}X^{\prime}$. There exists a canonical isomorphism of $X$-schemes
$\mathrm{Bl}_{\{Y_{i}^{\prime}:i\in I\}}^{\{D_{i}^{\prime}:i\in I\}}X^{\prime}\cong\mathrm{Bl}_{\{Y_{i}:i\in I\}}^{\{D_{i}:i\in I\}}X.$
@@
To simplify notation we set $X^{\prime\prime}=\mathrm{Bl}_{\{Y_{i}:i\in I\}}^{\{D_{i}:i\in I\}}X$. We have a canonical morphism of $X$-schemes $X^{\prime\prime}\to X^{\prime}$ which is an open immersion. (In fact this question is local on the target and explained at the beginning of the subsection in the local case). Setting $Y_{i}^{\prime\prime}=Y_{i}\times_{X}X^{\prime\prime}$ and $D_{i}^{\prime\prime}=D_{i}\times_{X}X^{\prime\prime}$ (for $i\in I$), by [Ma23b, §3.5] we therefore obtain canonical morphisms
$\mathrm{Bl}_{\{Y_{i}^{\prime\prime}:i\in I\}}^{\{D_{i}^{\prime\prime}:i\in I\}}X^{\prime\prime}\to\mathrm{Bl}_{\{Y_{i}^{\prime}:i\in I\}}^{\{D_{i}^{\prime}:i\in I\}}X^{\prime}\to\mathrm{Bl}_{\{Y_{i}:i\in I\}}^{\{D_{i}:i\in I\}}X.$
By [Ma23b, Proposition 3.28], we have $\mathrm{Bl}_{\{Y_{i}^{\prime\prime}:i\in I\}}^{\{D_{i}^{\prime\prime}:i\in I\}}X^{\prime\prime}=X^{\prime\prime}$. So we get $X$-morphisms $a:X^{\prime\prime}\to\mathrm{Bl}_{\{Y_{i}^{\prime}:i\in I\}}^{\{D_{i}^{\prime}:i\in I\}}X^{\prime}$ and $b:\mathrm{Bl}_{\{Y_{i}^{\prime}:i\in I\}}^{\{D_{i}^{\prime}:i\in I\}}X^{\prime}\to X^{\prime\prime}$. It is easily checked that $b\circ a=\mathrm{id}$ and $a\circ b=\mathrm{id}$ (e.g. using the universal property of dilatations), so these morphisms provide the desired identification. ∎
Proposition 3.29 ((Complementary)).
For all $i\in I$, $Y_{i}^{\prime}$ is a Cartier divisor on $X^{\prime}$. Moreover, the composition $\mathrm{Bl}^{\{Y_{i}^{\prime}:i\in I\}}_{\{\varnothing:i\in I\}}X^{\prime}\to X^{\prime}\to X$ is an open immersion; in fact, this is just the inclusion $X\smallsetminus(\cup_{i\in I}Y_{i})\to X$.
@@
First, assume that $X={\operatorname{Spec}}(A)$ for some ring $A$, and write $J_{i}$ for the global sections of ${\mathcal{J}}_{i}$. In this case, we have explained in the course of the proof of Proposition 3.27 that $\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$ is covered by affine open subschemes ${\operatorname{Spec}}\big{(}(\bigoplus_{\nu\in\mathbb{N}_{I}}J^{\nu})_{(\{a_{i}:i\in I\})}\big{)}$ where $(a_{i}:i\in I)$ is a family of elements of $A$ such that $a_{i}\in J_{i}$ for any $i$. As explained above, $(\bigoplus_{\nu\in\mathbb{N}_{I}}J^{\nu})_{(\{a_{i}:i\in I\})}$ is the dilatation ring $A[\{\frac{J_{i}}{a_{i}}:i\in I\}]$. Now the first assertion follows from [Ma23b, Facts 2.10 & 2.25]. The second assertion follows from the formula
(see e.g. the computation in [Ma23b, Proof of Corollary 2.23]) and the fact the the right-hand side is the localization of $A$ with respect to the $a_{i}$’s, see [Ma23b, Fact 2.11].
To treat the general case, one takes an affine open covering of $X$ and use the affine case. ∎
Proposition 3.30 ((Existence of blowups)).
Let $\pi:T\to X$ be a scheme such that, for all $i\in I$, $T\times_{X}Y_{i}$ is a Cartier divisor on $T$. Then there exists a unique morphism of $X$-schemes $T\to X^{\prime}$. As a consequence, the pair consisting of $X^{\prime}$ and the canonical morphism $X^{\prime}\to X$ is a blowup of $X$ with center $\{Y_{i}:i\in I\}$.
Proof.
It suffices to prove the statement in the affine case; namely we assume that $X={\operatorname{Spec}}(A)$ and $T={\operatorname{Spec}}(R)$ for some rings $A$ and $R$. Then we denote by $J_{i}$ the global sections of ${\mathcal{J}}_{i}$, so that $Y_{i}={\operatorname{Spec}}(A/J_{i})$. Then $\pi:T\to X$ corresponds to a morphism of rings $f:A\to R$. By assumption, for any $i\in I$, we have $f(J_{i})R=\gamma_{i}R$ for some element $\gamma_{i}\in R$ which is not a zero divisor.
Let us first show the existence of a morphism of $X$-schemes $T\to X^{\prime}$. We have a canonical morphism of $\mathbb{N}_{I}$-graded rings
$\bigoplus_{\nu\in\mathbb{N}_{I}}J^{\nu}\to\bigoplus_{\nu\in\mathbb{N}_{I}}(f(J)R)^{\nu}$
where $(f(J)R)^{\nu}$ denotes the ideal $\prod_{i\in I}(f(J_{i})R)^{\nu_{i}}\subset R$. Now since $f(J_{i})R=\gamma_{i}R$ with $\gamma_{i}$ non-zero divisor for any $i$, we have a canonical isomorphism of $R$-algebras $\bigoplus_{\nu\in\mathbb{N}_{I}}(f(J)R)^{\nu}\cong R[\mathbb{N}_{I}]$. So we get a canonical morphism of graded $A$-algebras
$\bigoplus_{\nu\in\mathbb{N}_{I}}J^{\nu}\to R[\mathbb{N}_{I}].$
By Example 3.19 we have ${\operatorname{Proj}}^{\mathbb{Z}_{I}}(R[\mathbb{N}_{I}])={\operatorname{Spec}}(R)$, so that by Proposition 3.12 this morphism provides a morphism of $X$-schemes $\phi:{\operatorname{Spec}}(R)\to{\operatorname{Proj}}(\bigoplus_{\nu\in\mathbb{N}_{I}}J^{\nu})$ as required.
@@
To show uniqueness we have to show that if $\phi^{\prime}:T\to X^{\prime}$ is an arbitrary morphism of $X$-schemes, we have $\phi=\phi^{\prime}$. By lemma 3.26, the morphism $X^{\prime}\to X$ is separated; as a consequence, the difference kernel $\mathrm{ker}(\phi,\phi^{\prime})$ (see [GW20, Definition 9.1]) is a closed subscheme of $T$ by [GW20, Definition/Proposition 9.7]. Let $U\subset T$ be the complement of the closed subscheme defined by the ideal generated by the $\gamma_{i}$’s; then the scheme theoretic image of the open immersion $\iota_{U}:U\to T$ is $T$ by [StP, Tag 056A]. Since $U\times_{X}Y_{i}=\varnothing$ for all $i$, the universal property of dilatations (see [Ma23b, Proposition 3.17]) implies that both $\phi_{|U}$ and $\phi^{\prime}_{|U}$ both factor through $\mathrm{Bl}_{\{\varnothing:i\in I\}}^{\{Y_{i}^{\prime}:i\in I\}}X^{\prime}$. By Proposition 3.29, $\mathrm{Bl}_{\{\varnothing:i\in I\}}^{\{Y_{i}^{\prime}:i\in I\}}X^{\prime}$ is an open subscheme of $X$, so that this implies that $\phi_{|U}=\phi^{\prime}_{|U}$. We deduce that $U\subset\mathrm{ker}(\phi,\phi^{\prime})$, and then that $\mathrm{ker}(\phi,\phi^{\prime})=T$, so that $\phi=\phi^{\prime}$.
@@
What we have now proved and Proposition 3.29 imply that $X^{\prime}$ satisfy the defining property of blowups (see Definition 3.25). ∎
@@
Recall the “$+$” operation on closed subschemes of $X$, see [Ma23b, Notation 3.1].
@@
We have a canonical isomorphism $\underline{{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X\cong\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$.
@@
It is enough to prove that $\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$ is also a blowup of $X$ with center $\{Y_{i}:i\in I\}$. This is immediate using that $\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$ is a blowup of $X$ with center $\{\sum_{i\in I}Y_{i}\}$ and the fact that Cartier divisors form a face of the monoid of closed subschemes (see [Ma23b, Notation 3.1]). ∎

### v4

@@ 3.6 Proj schemes, blowups and dilatations
In the setting above, the scheme ${\operatorname{Proj}}^{\mathbb{Z}^{I}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)$ is covered by the open subschemes $D_{\dagger}(\{a_{i}:i\in I\})$ where $\{a_{i}:i\in I\}$ runs over the families such that $a_{i}\in J_{i}$ for all $i\in I$.
@@
By Corollary 3.10, the lemma will follow if we prove that any non nilpotent relevant element in the $\mathbb{Z}^{I}$-graded ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$ belongs to the sum of the ideals $({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)\cdot\prod_{i}a_{i}$ where $\{a_{i}:i\in I\}$ runs over the families as above. Now it is clear that this sum is the sum of the graded components in ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$ associated with elements in $\mathbb{Z}^{I}$ all of whose components are positive. The claim therefore follows from the observation that the degree of any non nilpotent relevant element in ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$ must have positive components, because if the $i$-component of the degree of a non nilpotent homogeneous element is 00, then the same will hold for all divisors of all of its powers, so that this element cannot be relevant. ∎
We now work in a relative setting; we therefore fix a base scheme $X$, and let $\{Y_{i}:i\in I\}$ be a family of closed subschemes parametrized by a finite set $I$. We first recall the definition of blowups.
@@
Clearly, a blowup of $X$ with center $\{Y_{i}:i\in I\}$ is unique up to unique isomorphism if it exists. For each $i$, let $\mathcal{J}_{i}\subset{\mathcal{O}}_{X}$ be the quasi-coherent ideal which defines $Y_{i}$, and let
$\mathrm{Bl}_{\{\mathcal{J}_{i}:i\in I\}}\mathcal{O}_{X}=\bigoplus_{\nu\in\mathbb{N}^{I}}\mathcal{J}^{\nu}$
be the associated Rees algebra (cf. e.g. [Ma23b, §3.1]). This is canonically a $\mathbb{Z}^{I}$-graded quasi-coherent $\mathcal{O}_{X}$-algebra. Using the construction of §3.5 we now define an $X$-scheme
$\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X:={\operatorname{Proj}}^{\mathbb{Z}^{I}}_{X}(\mathrm{Bl}_{\{\mathcal{J}_{i}:i\in I\}}\mathcal{O}_{X}).$
We will prove in Proposition 3.28 below that this scheme is a blowup of $X$ with center $\{Y_{i}:i\in I\}$ in the sense of Definition 3.24. (The case $|I|=1$ of this statement is classical, see e.g. [StP, Tag 01OF], but our proof is different.)
@@
Since the property of being separated is local on the target (see e.g. [StP, Tag 02KU]), we can assume $X={\operatorname{Spec}}(A)$ is affine, and by [StP, Tag 01KV] the claim will follow if in this case we prove that $\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$ is a separated scheme. We proceed with the usual local notations from the beginning of the subsection. Consider the collection of families $\{a_{i}:i\in I\}$ with $a_{i}\in J_{i}$ for all $i\in I$, and let $\mathcal{F}\subset\mathcal{F}_{A}$ be the subset consisting of the multiplicative subsets generated by such families. By Lemma 3.23, we have ${\operatorname{Proj}}^{\mathbb{Z}^{I}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)={\operatorname{Proj}}^{\mathbb{Z}^{I}}_{{\mathcal{F}}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)$. Using the notation from Proposition 3.7, for any $T\in{\mathcal{F}}$ we have $C_{T}=(\mathbb{R}_{\geqslant 0})^{I}\subset\mathbb{R}^{I}$. So if $T,T^{\prime}$ belong to ${\mathcal{F}}$, $C_{T}\cap C_{T^{\prime}}$ has nonempty interior. Now Proposition 3.7 finishes the proof. ∎
@@
This statement is local on $X$, so we can assume $X={\operatorname{Spec}}(A)$ for some ring $A$, and write $J_{i}$ for the global sections of ${\mathcal{J}}_{i}$. In view of Lemma 3.23, in this case it suffices to prove that for any family $\{a_{i}:i\in I\}$ with $a_{i}\in J_{i}$ for any $i$, if $C$ is an effective Cartier divisor on $X$ then ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])\times_{X}C$ is an effective Cartier divisor on ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])$. This statement is a special case of [Ma23b, Proposition 3.18]. ∎
In the following statement we use the notation of dilatations of schemes from [Ma23b] (see also [MRR23, DMdS23]). From now on, for simplicity of notation we set
$X^{\prime}:=\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X,\quad Y_{i}^{\prime}:=Y_{i}\times_{X}X^{\prime}.$
Proposition 3.27 ((Complementary)).
For all $i\in I$, $Y_{i}^{\prime}$ is an effective Cartier divisor on $X^{\prime}$. Moreover, the composition $\mathrm{Bl}^{\{Y_{i}^{\prime}:i\in I\}}_{\{\varnothing:i\in I\}}X^{\prime}\to X^{\prime}\to X$ is an open immersion; in fact, this is just the inclusion $X\smallsetminus(\cup_{i\in I}Y_{i})\to X$.
@@
This statement is local on $X$, so we can assume $X={\operatorname{Spec}}(A)$ for some ring $A$, and write $J_{i}$ for the global sections of ${\mathcal{J}}_{i}$. In this case, by Lemma 3.23 the scheme $X^{\prime}=\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$ is covered by affine open subschemes ${\operatorname{Spec}}\big{(}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)_{(\{a_{i}:i\in I\})}\big{)}$ where $\{a_{i}:i\in I\}$ is a family of elements of $A$ such that $a_{i}\in J_{i}$ for any $i$. As explained above, $({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)_{(\{a_{i}:i\in I\})}$ is the dilatation ring $A[\{\frac{J_{i}}{a_{i}}:i\in I\}]$. Now the first assertion follows from [Ma23b, Facts 2.10 & 2.25]. The second assertion follows from the formula
(see e.g. the computation in [Ma23b, Proof of Corollary 2.23]) and the fact the the right-hand side is the localization of $A$ with respect to the $a_{i}$’s, see [Ma23b, Fact 2.11]. ∎
Proposition 3.28 ((Existence of blowups)).
Let $\pi:T\to X$ be a scheme such that, for all $i\in I$, $T\times_{X}Y_{i}$ is an effective Cartier divisor on $T$. Then there exists a unique morphism of $X$-schemes $T\to X^{\prime}$. As a consequence, the pair consisting of $X^{\prime}$ and the canonical morphism $X^{\prime}\to X$ is a blowup of $X$ with center $\{Y_{i}:i\in I\}$.
@@
First we show that for $T$ and $\pi$ as in the statement, there exists at most one morphism of $X$-schemes $T\to X^{\prime}$. In fact, let $\phi,\phi^{\prime}:T\to X^{\prime}$ be two morphisms of $X$-schemes. Denote by $U$ the open complement of the sum of the effective Cartier divisors $T\times_{X}Y_{i}$; then $U$ is scheme theoretically dense in $T$ by [StP, Tag 07ZU]. By Lemma 3.25, the morphism $X^{\prime}\to X$ is separated; as a consequence, the difference kernel $\mathrm{ker}(\phi,\phi^{\prime})$ (see [GW20, Definition 9.1]) is a closed subscheme of $T$ by [GW20, Definition/Proposition 9.7]. Since $U\times_{X}Y_{i}=\varnothing$ for all $i$, the universal property of dilatations (see [Ma23b, Proposition 3.17]) implies that both $\phi_{|U}$ and $\phi^{\prime}_{|U}$ factor through $\mathrm{Bl}_{\{\varnothing:i\in I\}}^{\{Y_{i}^{\prime}:i\in I\}}X^{\prime}$. By Proposition 3.27, $\mathrm{Bl}_{\{\varnothing:i\in I\}}^{\{Y_{i}^{\prime}:i\in I\}}X^{\prime}$ is an open subscheme of $X$, so that this implies that $\phi_{|U}=\phi^{\prime}_{|U}$. We deduce that $U\subset\mathrm{ker}(\phi,\phi^{\prime})$, and then that $\mathrm{ker}(\phi,\phi^{\prime})=T$, so that $\phi=\phi^{\prime}$.
To prove existence we first consider the following setting. We assume that $X={\operatorname{Spec}}(A)$ and $T={\operatorname{Spec}}(R)$ for some rings $A$ and $R$, and denote by $J_{i}$ the global sections of ${\mathcal{J}}_{i}$, so that $Y_{i}={\operatorname{Spec}}(A/J_{i})$, and by $f:A\to R$ the morphism of rings corresponding to $\pi:T\to X$. With this notation, we further assume that, for any $i\in I$, we have $f(J_{i})R=\gamma_{i}R$ for some non-zero-divisor $\gamma_{i}\in R$.
Let us first show the existence of a morphism of $X$-schemes $T\to X^{\prime}$. The morphism $f$ induces a canonical morphism of $\mathbb{Z}^{I}$-graded rings
$\bigoplus_{\nu\in\mathbb{N}^{I}}J^{\nu}\to\bigoplus_{\nu\in\mathbb{N}^{I}}(f(J)R)^{\nu}$
where $(f(J)R)^{\nu}$ denotes the ideal $\prod_{i\in I}(f(J_{i})R)^{\nu_{i}}\subset R$. Now since $f(J_{i})R=\gamma_{i}R$ with $\gamma_{i}$ non-zero-divisor for any $i$, we have a canonical isomorphism of $R$-algebras $\bigoplus_{\nu\in\mathbb{N}^{I}}(f(J)R)^{\nu}\cong R[\mathbb{N}^{I}]$. So we get a canonical morphism of graded $A$-algebras
$\bigoplus_{\nu\in\mathbb{N}^{I}}J^{\nu}\to R[\mathbb{N}^{I}].$
By Example 3.19 we have ${\operatorname{Proj}}^{\mathbb{Z}^{I}}(R[\mathbb{N}^{I}])={\operatorname{Spec}}(R)$, so that by Proposition 3.12 this morphism provides a morphism of $X$-schemes $\phi$ from an open subscheme of ${\operatorname{Spec}}(R)$ to ${\operatorname{Proj}}(\bigoplus_{\nu\in\mathbb{N}_{I}}J^{\nu})$. We will now prove that this open subscheme is the whole of ${\operatorname{Spec}}(R)$, which will conclude the proof of existence. In fact, ${\operatorname{Proj}}^{\mathbb{Z}^{I}}(R[\mathbb{N}^{I}])$ coincides with its open subscheme associated with the relevant element corresponding to the element in $\mathbb{N}^{I}$ all of whose components are $1$, which corresponds to $\prod_{i\in I}\gamma_{i}\in\bigoplus_{\nu\in\mathbb{N}^{I}}(f(J)R)^{\nu}$. Now for any $i$ there exist finite collections $(\delta_{i,j}:j\in A_{i})$ of elements of $J_{i}$ and $(r_{i,j}:j\in A_{i})$ of elements of $R$ such that $\gamma_{i}=\sum_{j\in A_{i}}r_{i,j}f(\delta_{i,j})$. Then we have
$\prod_{i\in I}\gamma_{i}=\sum_{(j_{i})\in\prod_{i}A_{i}}\prod_{i}r_{i,j_{i}}\cdot\prod_{i}f(\delta_{i,j_{i}}).$
In view of the construction of §3.3 and Lemma 3.5(1), this implies that ${\operatorname{Proj}}^{\mathbb{Z}^{I}}(R[\mathbb{N}^{I}])$ is covered by the open subschemes $D_{\dagger}(\prod_{i}f(\delta_{i,j_{i}}))$. Now by construction $\phi$ is defined on each of these subsets, which finishes the justification.
Now we prove the existence of a morphism of $X$-schemes $T\to X^{\prime}$ in general. For this we choose an affine open covering $X=\bigcup_{\alpha\in A}X_{\alpha}$. Then for any $\alpha\in A$ and $i\in I$, $(T\times_{X}X_{\alpha})\times_{X_{\alpha}}(Y_{i}\times_{X}X_{\alpha})$ is an effective Cartier divisor on $T\times_{X}X_{\alpha}$. Hence there exists an affine open covering $T\times_{X}X_{\alpha}=\bigcup_{\beta\in B_{\alpha}}T_{\alpha,\beta}$ such that each $T_{\alpha,\beta}\times_{X_{\alpha}}(Y_{i}\times_{X}X_{\alpha})$ is the closed subscheme defined by a non-zero-divisor in $\mathcal{O}(T_{\alpha,\beta})$, see e.g. [StP, Tag 01WS]. By the case treated above, for any $\alpha,\beta$ there exists a morphism of $X_{\alpha}$-schemes $T_{\alpha,\beta}\to X^{\prime}\times_{X}X_{\alpha}$. By uniqueness, these morphisms coincide on the intersections of the $T_{\alpha,\beta}$’s, hence they glue to a morphism of $X$-schemes $T\to X^{\prime}$, as desired.
@@
What we have now proved and Proposition 3.27 imply that $X^{\prime}$ satisfies the defining property of blowups (see Definition 3.24), so that the proof is complete. ∎
@@
We finish this subsection with some applications of Proposition 3.28. For the first one, recall the “$+$” operation on closed subschemes of $X$, see [Ma23b, Notation 3.1].
@@
We have a canonical isomorphism of $X$-schemes $\underline{{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X\cong\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$.
@@
It is enough to prove that $\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$ is also a blowup of $X$ with center $\{Y_{i}:i\in I\}$. This is immediate using that $\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$ is a blowup of $X$ with center $\{\sum_{i\in I}Y_{i}\}$ and the fact that effective Cartier divisors form a face of the monoid of closed subschemes (see [Ma23b, Notation 3.1]). ∎
Corollary 3.31.
Let $K\subset I$ be a subset and set $J=I\smallsetminus K$. We have a canonical isomorphism of $X$-schemes
$\underline{{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X\cong\underline{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X.$
Proof.
It is enough to prove that $\underline{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$ is also a blowup of $X$ with center $\{Y_{i}:i\in I\}$. Using Proposition 3.26, we see that $\underline{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$ belongs to ${\operatorname{Sch}}_{X}^{\{Y_{i}:i\in I\}\text{-reg}}$. Let $T\to X$ be in ${\operatorname{Sch}}_{X}^{\{Y_{i}:i\in I\}\text{-reg}}$. Applying twice Proposition 3.28, we get a unique $X$-morphism $T\to\underline{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$. This finishes the proof. ∎